OpenStudy (anonymous):

Find the x -intercept(s) and the coordinates of the vertex for the parabola Y= x^2 + 6x - 7 . If there is more than one x-intercept, separate them with commas.

OpenStudy (anonymous):

So basically I find the vertex using the equation and then plot two points on either side of the vertex and then just find the x intercept of each point? @

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, the x-intercepts is where the graph touches the x-axis at that point, y = 0 so if you set that to 0, and solve "x", by factoring you should be able to get the "solutions" \(\bf 0 = x^2 + 6x - 7\) unless you're asked to graph it

OpenStudy (jdoe0001):

and yes, you could do it like the previous one as well get the vertex, a couple of points on either side, to get the x-intercepts or solutions

OpenStudy (jdoe0001):

\(\bf y= x^2 + 6x - 7 \\ \quad \\ \textit{vertex of a parabola}\\ \quad \\ y = {\color{red}{ 1}}x^2{\color{blue}{ +6}}x{\color{green}{ -7}}\qquad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (anonymous):

ok cool so I guess my only question now is how do you know how the equation plugs in to the other one like that?

OpenStudy (jdoe0001):

hmmm what do you mean?

OpenStudy (anonymous):

like when you color code the equation I post so I know where to plug it in at, how do you know that's where they plug in at? oh and I got (3,-2) btw

OpenStudy (jdoe0001):

ohhh is a notation, or formula based on the pattern of the values

OpenStudy (jdoe0001):

well \(\bf \left(-\cfrac{{\color{blue}{ 6}}}{2{\color{red}{ (1)}}}\quad ,\quad {\color{green}{ -7}}-\cfrac{{\color{blue}{ 6}}^2}{4{\color{red}{ (1)}}}\right)\implies \left( -3,-7-\cfrac{\cancel{36}}{\cancel{4}} \right) \\ \quad \\ \left( -3, -7-9 \right)\)

OpenStudy (anonymous):

ohh yeahh nope definitely got that wrong.. thank you though!

OpenStudy (jdoe0001):

yw